# Existence of Decreasing Nambu Solutions to the Rainbow–Ladder Gap Equation of QCD by Cone Compression

### v1

@@ Appendix C Tail of the quark mass function
Let’s extract the large-momentum behavior of the quark mass function. The gap equation in Landau gauge reads
[…]
with
$\Sigma(p)=\frac{4}{3}\int\frac{d^{4}q}{(2\pi)^{4}}4\pi\alpha_{s}(q)\gamma_{\mu}\Delta(q)P^{\mu\nu}(q)D(k)\Gamma_{\nu}(q,k,p)$ (20)
We perform the Wick rotation to Euclidean space and, in the large momentum limit, we can take $\Gamma_{\nu}(q,k,p)=\gamma_{\nu}$. Also, the logarithmic pole of the angular integral does not warrant including the gluon mass whose contribution is suppressed by $\frac{M_{g}^{2}}{p^{2}}$ at large $p$. We can borrow a result of the angular integral
@@ Appendix E Tail of the quark field renormalization function
Taking instead the trace of $\not${p} times the gap equation and with $\frac{1}{Z(p)}-1\equiv\tilde{A}(p)$, we obtain
$\tilde{A}(p)=\frac{1}{12\pi^{4}p^{2}}\int_{0}^{\infty}dkk^{3}\frac{\frac{1}{\tilde{A}(k)+1}}{k^{2}+M^{2}(k)}\int_{S^{4}}(p\cdot k+2p\cdot qk\cdot q/q^{2})\mathcal{G}(q^{2})$ (32)
We have
@@
with $\chi_{Z}(k,p)$ polynomially suppressed over the other term at large $k\gg p$. Therefore, for $\tilde{A}$ converging to 0 in the UV as is required by the renormalization condition at $\infty$, the renormalization scale can be taken to infinity in the gap equation. However, after converting to differential form as in Eqn.24, because of the $x^{2}$ factor, $\frac{1}{x^{n}}$ perturbations grow to $\frac{1}{x^{n-1}}$ in one iteration, making the impact of $\chi_{Z}(k,p)$ unpredictable. We prove in the main text that there exist $A_{1,2}$ such that if $A_{1}(p)<A(p)<A_{2}(p)$, then the same is true for $T(A)$ which circumvents this issue.

### v2

@@ Appendix C Tail of the quark mass function
Let’s extract the large-momentum behavior of the quark mass function. In this limit, the dominant contribution comes from $\mathcal{G}(q)\sim\frac{4\pi^{2}\gamma_{m}}{q^{2}\log(\frac{q^{2}}{\mu^{2}})}$.
We can borrow a result of the angular integral
@@ Appendix E Tail of the quark field renormalization function
Following Eqn.20 with $\mathcal{G}(q)\sim\frac{4\pi^{2}\gamma_{m}}{q^{2}\log(\frac{q^{2}}{\mu^{2}})}$ as the dominant contribution, we observe that
@@
with $\chi_{Z}(k,p)$ polynomially suppressed over the other term at large $k\gg p$. Therefore, for $A(p)$ converging to $1$ in the UV as required by the renormalization condition at $\infty$, the integral converges due to the square of the log in the denominator and the renormalization scale can be taken to infinity in the gap equation. However, after converting to differential form as in Eqn.24, because of the $x^{2}$ factor, $\frac{1}{x^{n}}$ perturbations grow to $\frac{1}{x^{n-1}}$ in one iteration, making the impact of $\chi_{Z}(k,p)$ unpredictable. We prove in the main text that there exist $A_{1,2}$ such that if $A_{1}(p)<A(p)<A_{2}(p)$, then the same is true for $T(A)$ which circumvents this issue.